# Optimal quantum simulation of linear non-unitary dynamics

### v1

@@ 3 Optimal scaling LCHS
In this section, we prove our main algorithmic result Theorem 2 that $\hat{f}_{2}(k;\gamma,c)$ leads to a LCHS that approximates $U_{0}(t)$ to error $\epsilon$ in spectral norm with the desired scaling $\gamma=\mathcal{O}(\frac{1}{c}\log^{1/2}\frac{1}{\epsilon})$ and $R=\mathcal{O}(\frac{1}{c}\log\frac{1}{\epsilon})$ for any $c>0$. We start by evaluating some useful properties of the entire functions $f_{2,1}$ Equation 9, summarized in Table 2. These functions satisfy
@@
Lemma 4 $f_{1}(x;\gamma)$ $\frac{1}{\sqrt{2\pi}}\frac{1}{1-ik}e^{-\frac{k^{2}+1}{4\gamma^{2}}}$ $\mathcal{O}(1)$ $\mathcal{O}(\log\gamma)$ $\mathcal{O}(\log\frac{1}{\epsilon})$
Lemma 5 $f_{2}(x;\gamma)$ $\sqrt{\frac{2}{\pi}}\frac{1}{1+k^{2}}e^{-\frac{k^{2}+1}{4\gamma^{2}}}$ $\text{erfc}\left(\frac{1}{2\gamma}\right)$ $\text{erfc}\left(\frac{1}{2\gamma}\right)$ $\mathcal{O}(\log\frac{1}{\epsilon})$
Corollary 6 $f_{2}(x;\gamma,c)$ $\sqrt{\frac{2}{\pi}}\frac{e^{c(1-ik)}}{1+k^{2}}e^{-\frac{k^{2}+1}{4\gamma^{2}}}$ $e^{c}\text{erfc}\left(\frac{1}{2\gamma}\right)$ $e^{c}\text{erfc}\left(\frac{1}{2\gamma}\right)$ $\mathcal{O}(\frac{1}{c}\log\frac{1}{\epsilon})$
Table 2: Examples of kernel functions whose inverse Fourier transform satisfies: (First two rows) Exact exponential decay Equation 4; (Last four rows) Approximate exponential decay Equation 5. $R>0$ is the truncation radius that leads to at most an error $\epsilon$ in the truncated integral $\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}\backslash[-R,R]}|\hat{f}(k)|\mathrm{d}k\leq\epsilon$. The approximate cases are parameterized by $R=2c\gamma^{2}=\Theta(\frac{1}{c}\log\frac{1}{\epsilon})$.
@@
and the proof is completed similar to Equation 44.
@@
and Lemma 4 for the Fourier transform of $\hat{f}(k;\gamma)$. The integral of the absolute value
@@
Equation 49 is proven by summing the upper bounds on $f_{1}$ Equations 41 and 42 according to the definition of $f_{2}$. By the convolution theorem, $f_{2}$ is proportional to the convolution of $e^{-|x|}$ and a Gaussian $e^{-x^{2}\gamma^{2}}$. As both functions have maximum at $x=0$, are symmetric, and have monotonic decay, $f_{2}$ is maximized at $x=0$. Together with Equation 52, this proves Equation 50. ∎
The approximations Lemmas 4 and 5 approximate $e^{-x}$ for $x\geq c$. An approximation for $x\geq 0$ is thus obtained by a translation and scale by $c$ and $e^{c}$ respectively.
@@
Proof of Theorem 2.
According to Theorem 1, the error of the LCHS
@@
In Lemma 7, we bound the integral
@@
In Lemma 8, we bound the integral
@@
Then we may upper bound Equation 77 by $\epsilon_{1}+\epsilon_{2}=\epsilon$ with the choice $r=\frac{1}{2\pi}$, $\epsilon_{1}=\frac{1}{1+r}\epsilon$ and $\epsilon_{2}=\frac{r}{1+r}\epsilon$. Then $\forall c>0$, we choose
@@
The bound $\alpha_{\hat{f},R}\leq\alpha_{\hat{f},\infty}\doteq\alpha_{\hat{f}}$ is true for any $\hat{f}$ as the integrand $|\hat{f}(k)|$ is positive, and in Corollary 6, we found that $\alpha_{\hat{f}_{2}}=\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}|\hat{f}_{2}(k;\gamma,c)|\mathrm{d}k=e^{c}\mathrm{erfc}\left(\frac{1}{2\gamma}\right)$, and for any $x\geq 0$, $\mathrm{erfc}(x)\leq 1$. ∎
@@
We choose $y_{0}=2c\gamma^{2}$, as it maximizes the exponential decay rate. With this choice, $\left|e^{-\frac{z^{2}+1}{4\gamma^{2}}-icz}\right|=e^{-c^{2}\gamma^{2}+-\frac{x^{2}+1}{4\gamma^{2}}}$. Using the assumption that $c\gamma^{2}\geq 1$, the integral is bounded by
@@
where we evaluate the integral with the identity $\mathrm{erfc}(a)=\frac{1}{\pi}\int_{\mathbb{R}}\frac{e^{-a^{2}(1+x^{2})}}{1+x^{2}}\mathrm{d}x$ for $a\in\mathbb{R}$ [OOL+, Equation.7.7.1] and we solve for the $\gamma$ that achieves the desired $\epsilon$. ∎
@@ 4 Exponentially convergent uniform quadrature
A gate-efficient implementation of LCHS requires discretizing the integral of $O_{R,\gamma,c}(t)$ to a finite number of quadrature points in a finite domain [R,R][-R,R]. In this section, we prove Theorem 3 that a uniform stepsize $h>0$, such that $R/h$ is an integer, with uniform weights
@@
$\left\|I_{R,\gamma,c,h}(t)-U_{0}(t)\right\|\leq\epsilon_{\text{lchs}}+\mathcal{O}\left(e^{\frac{1}{2}\|L\|_{L^{1}}-\pi/h}\right)\leq\epsilon_{\text{lchs}}+\epsilon_{\text{qaud}},$ (73)
where $\epsilon_{\text{lchs}}$ is the additive contribution from Theorem 2, and $h$ is chosen to achieve the desired error $\epsilon_{\text{quad}}$. We prove this using the well-known result [TW14] of an exponentially convergent trapezoid rule that we restate in Lemma 9, which applies to summands that are analytic on a finite strip in the complex plane. Note that Lemma 9 was originally proven for scalar $f$. However, the underlying techniques (e.g. the Poisson summation formula and contour integration), generalize to functions valued in a Banach space and so the absolute value $|f|$ generalizes to a matrix norm of $\|f\|$.
@@
We may apply Lemma 9 to Equation 72 as we have previously shown in Lemma 2 that $\mathcal{U}(t;z)$ is an entire function with respect to $z$, and $\hat{f}_{2}(z;\gamma,c)$ has poles only at $k=\pm i$.
Proof of Theorem 3.
@@
where the last line follows from Lemma 10, which proves that
@@
for any $\gamma\geq 0$ and integer $R/h>0$, and Theorem 1, which proves the error bound on
@@
for any $y_{0}>1$. In the proof of Theorem 2, we bounded the integrals in Lemma 7 and Lemma 8 to show that the last two terms of Equation 77 $\frac{1}{\sqrt{2\pi}}\left(\int_{I_{R}}|\hat{f}_{2}(k;\gamma,c)|\mathrm{d}k+\int_{\mathbb{R}}|\hat{f}_{2}(k-iy_{0};\gamma,c)|\mathrm{d}k\right)\leq\epsilon_{\text{lchs}}$ for any $\epsilon_{\text{lchs}}\in(0,1]$ and any $c>0$ if $R=2c\gamma^{2}$ and $\gamma=\frac{1}{c}\sqrt{c+\log\frac{1+1/({2\pi})}{\epsilon_{\text{lchs}}}}$. In Lemma 11, we prove that the uniform discretization Equation 72 contributes an error
@@
which is bounded by any $\epsilon_{\text{quad}}\in\left(0,\frac{4}{15}\right]$ for any $c>0$ with the any choice $h\leq\frac{\pi}{\|L\|_{L^{1}}/2+\log\left(\frac{64e^{3c/2}}{15\epsilon_{\text{quad}}}\right)}$. In Lemma 12, we prove that under the same parameter choices of $\gamma,R,h$, the block-encoding normalization
@@
where the last quantity is bounded in Lemma 8.
@@
In Lemma 2, we proved that $\mathcal{U}(t;z)$ is an entire function with respect to $z$ if $\int_{0}^{t}\|L(s)\|\mathrm{d}s<\infty$ and $\int_{0}^{t}\|H(s)\|\mathrm{d}s<\infty$. Hence, $o(z)$ has poles only at $\pm i$. In other words, $o(z)$ is analytic on the strip $S_{1}$. We apply the following inequalities to each term in $o(z)$, which hold for any $z=x+iy\in S_{1}$}.
@@
These imply uniform convergence of $\lim_{x\rightarrow\pm\infty}o(x+iy)=0$ for any finite $y$. We now evaluate the error pre-factor $M$ of Lemma 9. For any $y\in(-1,1)$,
@@
where the last equality follows from the identity $\mathrm{erfc}(a)=\frac{1}{\pi}e^{-a^{2}}\int_{\mathbb{R}}\frac{e^{-a^{2}x^{2}}}{1+x^{2}}$ [OOL+, Equation.7.7.1]. Hence the maximum over $y\in(-a,a)$ for $a\in(0,1)$ is bounded by
@@
Now apply Lemma 9 to obtain the quadrature error
@@
By Lemma 10 in the case $\mathcal{U}(t;k)=1$, the first term
@@
By Lemma 8, this is further bounded by
@@
By Lemma 11 in the case $H(s)=L(s)=0$, the second term
@@ 5 Gate-efficient circuit implementation
In this section, we prove our main result on an explicit LCHS quantum circuit implementation Theorem 4 for block-encoding
@@
A standard block-encoding construction is linear-combination-of-unitaries Lemma 13. In prior art [ACL23], it was already known that the query complexity of this block-encoding is no more than that of Hamiltonian simulation of $\mathcal{U}(t;\pm R)$. For completeness, we review this construction, which is used to synthesize the Sel unitary, and then provide an explicit circuit construction of Prep for our kernel function $\hat{f}_{2}$.
@@
In the query setting, the time-ordered evolution operator is constructed from either the block-encoding $\textsc{Be}[A/\alpha_{A}]$ in the time-independent case or the time-dependent block-encoding HamT Definition 14. Despite the additional controls by $j$, the query complexity of Sel Equation 110 is no more than that for $U_{0}(t)$.
@@
By linear-combination-of-unitaries, Lemma 13, these block-encode $\textsc{Be}\left[\frac{I_{h}}{|\vec{a}|_{1}}\right]$. By definition, note that $|\vec{a}|_{1}=\alpha_{\hat{f},R,h}$, and by Theorem 3, $\alpha_{\hat{f}_{2},R,h}=\alpha_{\hat{f}_{2}}+\mathcal{O}(\epsilon)$. The small difference between $\alpha_{\hat{f}_{2},R,h}$ and $\alpha_{\hat{f}_{2}}$ only changes the block-encoding normalization of $I_{h}$ by a small factor. We use the following lemmas to prove the claims gate and query complexity of this block-encoding.
  • In Lemma 18, we prove that the unitaries $\textsc{Prep}_{\lvert\vec{a}\rangle}$ and $\overline{\textsc{Prep}_{\lvert\vec{a}\rangle}}$ may be approximated by a unitary quantum circuit to error $\epsilon_{2}$ for any $\epsilon_{2}>0$, any $c>0$, any $h>0$, any $\gamma>0$, and any $R=\Omega(\gamma\log^{1/2}\frac{1}{\epsilon_{2}})$, using $G=\mathcal{O}(R^{3/2}\log\frac{R}{\epsilon_{2}}\log\frac{R}{h})$ two-qubit gates.
  • In Lemma 15, we prove that the query complexity $Q$ and gate complexity of Sel is no more than that of $\mathcal{U}(t;hj)$, plus $\mathcal{O}(Q\log\frac{R}{h})$ two-qubit gates
Let us choose all errors $\epsilon_{j}=\mathcal{O}(\epsilon)$. In Theorems 2 and 3, we chose $\gamma=\Omega(\frac{1}{c}\log^{1/2}\frac{1}{\epsilon_{\text{lchs}}})$, $R=\Omega(\frac{1}{c}\log\frac{1}{\epsilon_{\text{lchs}}})$, and $h^{-1}=\mathcal{O}(\|L\|_{L^{1}}+c+\log\frac{1}{\epsilon_{\text{quad}}})$. Assume that $c$ is a constant. Then under the above conditions, the total gate complexity is

### v2

@@ 3 Optimal scaling LCHS
In this section, we prove our main algorithmic result Theorem˜2 that $\hat{f}_{2}(k;\gamma,c)$ leads to a LCHS that approximates $U_{0}(t)$ to error $\epsilon$ in spectral norm with the desired scaling $\gamma=\mathcal{O}(\frac{1}{c}\log^{1/2}\frac{1}{\epsilon})$ and $R=\mathcal{O}(\frac{1}{c}\log\frac{1}{\epsilon})$ for any $c>0$. We start by evaluating some useful properties of the entire functions $f_{2,1}$ Equation˜9, summarized in Table˜2. These functions satisfy
@@
Lemma˜5 $f_{1}(x;\gamma)$ $\frac{1}{\sqrt{2\pi}}\frac{1}{1-ik}e^{-\frac{k^{2}+1}{4\gamma^{2}}}$ $\mathcal{O}(1)$ $\mathcal{O}(\log\gamma)$ $\mathcal{O}(\log\frac{1}{\epsilon})$
Lemma˜6 $f_{2}(x;\gamma)$ $\sqrt{\frac{2}{\pi}}\frac{1}{1+k^{2}}e^{-\frac{k^{2}+1}{4\gamma^{2}}}$ $\text{erfc}\left(\frac{1}{2\gamma}\right)$ $\text{erfc}\left(\frac{1}{2\gamma}\right)$ $\mathcal{O}(\log\frac{1}{\epsilon})$
Corollary˜7 $f_{2}(x;\gamma,c)$ $\sqrt{\frac{2}{\pi}}\frac{e^{c(1-ik)}}{1+k^{2}}e^{-\frac{k^{2}+1}{4\gamma^{2}}}$ $e^{c}\text{erfc}\left(\frac{1}{2\gamma}\right)$ $e^{c}\text{erfc}\left(\frac{1}{2\gamma}\right)$ $\mathcal{O}(\frac{1}{c}\log\frac{1}{\epsilon})$
Table 2: Examples of kernel functions whose inverse Fourier transform satisfies: (First two rows) Exact exponential decay Equation˜4; (Last four rows) Approximate exponential decay Equation˜5. $R>0$ is the truncation radius that leads to at most an error $\epsilon$ in the truncated integral $\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}\backslash[-R,R]}|\hat{f}(k)|\mathrm{d}k\leq\epsilon$. The approximate cases are parameterized by $R=2c\gamma^{2}=\Theta(\frac{1}{c}\log\frac{1}{\epsilon})$.
@@
and the proof is completed similar to Equation˜47.
@@
and Lemma˜5 for the Fourier transform of $\hat{f}(k;\gamma)$. The integral of the absolute value
@@
Equation˜52 is proven by summing the upper bounds on $f_{1}$ Equations˜44 and 45 according to the definition of $f_{2}$. By the convolution theorem, $f_{2}$ is proportional to the convolution of $e^{-|x|}$ and a Gaussian $e^{-x^{2}\gamma^{2}}$. As both functions have maximum at $x=0$, are symmetric, and have monotonic decay, $f_{2}$ is maximized at $x=0$. Together with Equation˜55, this proves Equation˜53. ∎
The approximations Lemmas˜5 and 6 approximate $e^{-x}$ for $x\geq c$. An approximation for $x\geq 0$ is thus obtained by a translation and scale by $c$ and $e^{c}$ respectively.
@@
Proof of Theorem˜2.
According to Theorem˜1, the error of the LCHS
@@
In Lemma˜8, we bound the integral
@@
In Lemma˜9, we bound the integral
@@
Then we may upper bound Equation˜60 by $\epsilon_{1}+\epsilon_{2}=\epsilon$ with the choice $r=\frac{1}{2\pi}$, $\epsilon_{1}=\frac{1}{1+r}\epsilon$ and $\epsilon_{2}=\frac{r}{1+r}\epsilon$. Then $\forall c>0$, we choose
@@
The bound $\alpha_{\hat{f},R}\leq\alpha_{\hat{f},\infty}\doteq\alpha_{\hat{f}}$ is true for any $\hat{f}$ as the integrand $|\hat{f}(k)|$ is positive, and in Corollary˜7, we found that $\alpha_{\hat{f}_{2}}=\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}|\hat{f}_{2}(k;\gamma,c)|\mathrm{d}k=e^{c}\mathrm{erfc}\left(\frac{1}{2\gamma}\right)$, and for any $x\geq 0$, $\mathrm{erfc}(x)\leq 1$. ∎
@@
We choose $y_{0}=2c\gamma^{2}$, as it maximizes the exponential decay rate. With this choice, $\left|e^{-\frac{z^{2}+1}{4\gamma^{2}}-icz}\right|=e^{-c^{2}\gamma^{2}-\frac{x^{2}+1}{4\gamma^{2}}}$. Using the assumption that $c\gamma^{2}\geq 1$, the integral is bounded by
@@
where we evaluate the integral with the identity $\mathrm{erfc}(a)=\frac{1}{\pi}\int_{\mathbb{R}}\frac{e^{-a^{2}(1+x^{2})}}{1+x^{2}}\mathrm{d}x$ for $a\in\mathbb{R}$ [OOL+25, Equation.7.7.1] and we solve for the $\gamma$ that achieves the desired $\epsilon$. ∎
@@ 4 Exponentially convergent uniform quadrature
A gate-efficient implementation of LCHS requires discretizing the integral of $O_{R,\gamma,c}(t)$ to a finite number of quadrature points in a finite domain [R,R][-R,R]. In this section, we prove Theorem˜3 that a uniform stepsize $h>0$, such that $R/h$ is an integer, with uniform weights
@@
$\left\|I_{R,\gamma,c,h}(t)-U_{0}(t)\right\|\leq\epsilon_{\text{lchs}}+\mathcal{O}\left(e^{\frac{1}{2}\|L\|_{L^{1}}-\pi/h}\right)\leq\epsilon_{\text{lchs}}+\epsilon_{\text{quad}},$ (76)
where $\epsilon_{\text{lchs}}$ is the additive contribution from Theorem˜2, and $h$ is chosen to achieve the desired error $\epsilon_{\text{quad}}$. We prove this using the well-known result [TW14] of an exponentially convergent trapezoid rule that we restate in Lemma˜10, which applies to summands that are analytic on a finite strip in the complex plane. Note that Lemma˜10 was originally proven for scalar $f$. However, the underlying techniques (e.g. the Poisson summation formula and contour integration), generalize to functions valued in a Banach space and so the absolute value $|f|$ generalizes to a matrix norm of $\|f\|$.
@@
We may apply Lemma˜10 to Equation˜75 as we have previously shown in Lemma˜3 that $\mathcal{U}(t;z)$ is an entire function with respect to $z$, and $\hat{f}_{2}(z;\gamma,c)$ has poles only at $k=\pm i$.
Proof of Theorem˜3.
@@
where the last line follows from Lemma˜11, which proves that
@@
for any $\gamma\geq 0$ and integer $R/h>0$, and Theorem˜1, which proves the error bound on
@@
for any $y_{0}>1$. In the proof of Theorem˜2, we bounded the integrals in Lemma˜8 and Lemma˜9 to show that the last two terms of Equation˜80 $\frac{1}{\sqrt{2\pi}}\left(\int_{I_{R}}|\hat{f}_{2}(k;\gamma,c)|\mathrm{d}k+\int_{\mathbb{R}}|\hat{f}_{2}(k-iy_{0};\gamma,c)|\mathrm{d}k\right)\leq\epsilon_{\text{lchs}}$ for any $\epsilon_{\text{lchs}}\in(0,1]$ and any $c>0$ if $R=2c\gamma^{2}$ and $\gamma=\frac{1}{c}\sqrt{c+\log\frac{1+1/({2\pi})}{\epsilon_{\text{lchs}}}}$. In Lemma˜12, we prove that the uniform discretization Equation˜75 contributes an error
@@
which is bounded by any $\epsilon_{\text{quad}}\in\left(0,\frac{4}{15}\right]$ for any $c>0$ with the any choice $h\leq\frac{\pi}{\|L\|_{L^{1}}/2+\log\left(\frac{64e^{3c/2}}{15\epsilon_{\text{quad}}}\right)}$. In Lemma˜13, we prove that under the same parameter choices of $\gamma,R,h$, the block-encoding normalization
@@
where the last quantity is bounded in Lemma˜9.
@@
In Lemma˜3, we proved that $\mathcal{U}(t;z)$ is an entire function with respect to $z$ if $\int_{0}^{t}\|L(s)\|\mathrm{d}s<\infty$ and $\int_{0}^{t}\|H(s)\|\mathrm{d}s<\infty$. Hence, $o(z)$ has poles only at $\pm i$. In other words, $o(z)$ is analytic on the strip $S_{1}$. We apply the following inequalities to each term in $o(z)$, which hold for any $z=x+iy\in S_{1}$}.
@@
These imply uniform convergence of $\lim_{x\rightarrow\pm\infty}o(x+iy)=0$ for any finite $y$. We now evaluate the error pre-factor $M$ of Lemma˜10. For any $y\in(-1,1)$,
@@
where the last equality follows from the identity $\mathrm{erfc}(a)=\frac{1}{\pi}e^{-a^{2}}\int_{\mathbb{R}}\frac{e^{-a^{2}x^{2}}}{1+x^{2}}$ [OOL+25, Equation.7.7.1]. Hence the maximum over $y\in(-a,a)$ for $a\in(0,1)$ is bounded by
@@
Now apply Lemma˜10 to obtain the quadrature error
@@
By Lemma˜11 in the case $\mathcal{U}(t;k)=1$, the first term
@@
By Lemma˜9, this is further bounded by
@@
By Lemma˜12 in the case $H(s)=L(s)=0$, the second term
@@ 5 Gate-efficient circuit implementation
In this section, we prove our main result on an explicit LCHS quantum circuit implementation Theorem˜4 for block-encoding
@@
A standard block-encoding construction is linear-combination-of-unitaries Lemma˜14. In prior art [ACL23], it was already known that the query complexity of this block-encoding is no more than that of Hamiltonian simulation of $\mathcal{U}(t;\pm R)$. For completeness, we review this construction, which is used to synthesize the Sel unitary, and then provide an explicit circuit construction of Prep for our kernel function $\hat{f}_{2}$.
@@
In the query setting, the time-ordered evolution operator is constructed from either the block-encoding $\textsc{Be}[A/\alpha_{A}]$ in the time-independent case or the time-dependent block-encoding HamT Definition˜15. Despite the additional controls by $j$, the query complexity of Sel Equation˜113 is no more than that for $U_{0}(t)$.
@@
By linear-combination-of-unitaries, Lemma˜14, these block-encode $\textsc{Be}\left[\frac{I_{h}}{|\vec{a}|_{1}}\right]$. By definition, note that $|\vec{a}|_{1}=\alpha_{\hat{f},R,h}$, and by Theorem˜3, $\alpha_{\hat{f}_{2},R,h}=\alpha_{\hat{f}_{2}}+\mathcal{O}(\epsilon)$. The small difference between $\alpha_{\hat{f}_{2},R,h}$ and $\alpha_{\hat{f}_{2}}$ only changes the block-encoding normalization of $I_{h}$ by a small factor. We use the following lemmas to prove the claims gate and query complexity of this block-encoding.
  • In Lemma˜19, we prove that the unitaries $\textsc{Prep}_{\lvert\vec{a}\rangle}$ and $\overline{\textsc{Prep}_{\lvert\vec{a}\rangle}}$ may be approximated by a unitary quantum circuit to error $\epsilon_{2}$ for any $\epsilon_{2}>0$, any $c>0$, any $h>0$, any $\gamma>0$, and any $R=\Omega(\gamma\log^{1/2}\frac{1}{\epsilon_{2}})$, using $G=\mathcal{O}(R^{3/2}\log\frac{R}{\epsilon_{2}}\log\frac{R}{h})$ two-qubit gates.
  • In Lemma˜16, we prove that the query complexity $Q$ and gate complexity of Sel is no more than that of $\mathcal{U}(t;hj)$, plus $\mathcal{O}(Q\log\frac{R}{h})$ two-qubit gates
Let us choose all errors $\epsilon_{j}=\mathcal{O}(\epsilon)$. In Theorems˜2 and 3, we chose $\gamma=\Omega(\frac{1}{c}\log^{1/2}\frac{1}{\epsilon_{\text{lchs}}})$, $R=\Omega(\frac{1}{c}\log\frac{1}{\epsilon_{\text{lchs}}})$, and $h^{-1}=\mathcal{O}(\|L\|_{L^{1}}+c+\log\frac{1}{\epsilon_{\text{quad}}})$. Assume that $c$ is a constant. Then under the above conditions, the total gate complexity is